# Libra: Leveraging Temporal Images for Biomedical Radiology Analysis

### v1

@@ 3.1 Task and Dataset
Inspired by the multi-stage training techniques commonly employed in recent MLLMs fine-tuning methods (Zhu et al., 2023), we adopt a two-stage training strategy to infuse multimodal capabilities into the model. In the first stage, we leverage a large-scale dataset to train the adapter, primarily focusing on learning the distribution of CXR token embeddings and the general relationships between CXR and text tokens across the MIMIC dataset. The second stage fine-tuning the model on downstream tasks, where the text decoder (i.e., the LLM) is refined to focus on aligning high-granularity information between CXRs and the Findings section of radiology reports.
In addition to Findings section generation, the temporal feature alignment stage also uses the Impression section and VQA tasks. The Impression section summarises diagnoses and proposes further investigations (Babar et al., 2021), aiding the alignment between CXRs and their textual descriptions. We use the same system and clinical prompts as for Findings, replacing ‘Findings’ with ‘Impression’. For VQA, system prompts remain the same, while clinical prompts adapt to medical-specific questions, guiding the model to generate corresponding captions. These fine-grained VQA tasks expand the MLLM’s biomedical vocabulary and improve the alignment of image-text annotations.
3.2 Evaluation Metrics
We evaluate the generated reports using lexical and radiology-specific metrics, following established protocols.. For lexical metrics, we use ROUGE-L (Lin, 2004), BLEU- (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005) and BERT (Devlin et al., 2019). For radiology-specific metrics, we use RadGraph-F1 (Jain et al., 2021), RG (Delbrouck et al., 2022a), F1-CheXpert (Irvin et al., 2019a), CheXbert vector similarity (Yu et al., 2022), as well as RadCliQ version Yu et al. (2022).
Additionally, we introduce a novel temporal entity F1 score (called ) to assess the model’s ability to capture temporal information. Clinical metrics typically emphasise the accuracy of medical findings, prioritising the detection of clinically relevant entities. In contrast, the temporal entity score specifically measures the accuracy of entities related to progression over time described in the report. Detailed descriptions of all metrics and an illustrative analysis are provided in Appx. B.3.
Temporal Entity F1 Following the work of Bannur et al. (2023), we set a reward list that includes common radiology-related keywords associated with temporal changes. Temporal entities are then extracted from both the ground truth and the generated reports . Importantly, no stemming or lemmatization is applied during token processing to preserve the precise description of temporal changes. After extraction, we compute precision and recall , which are subsequently used to calculate the temporal entity score, defined as the harmonic mean of precision and recall (Van Rijsbergen, 1974), also known as the F1 score.
where is a small value, set to a default of , to prevent division by zero (it is also added to the numerator for special cases where no temporal entities are present in the ground truth). The coefficient controls the balance between precision and recall, with a default value of .
3.3 Main Result
Although MIMIC-CXR provides an “official” test split, strict comparisons with prior studies remain challenging due to differences in test set inclusion criteria and pre-processing steps. For instance, Yu et al. (2022) and Jeong et al. (2023) included only one image per study, resulting in a test set of 1,597 samples, while Tanida et al. (2023) followed the Chest ImaGenome split (Wu et al., 2021b). Such variations in test set distributions can significantly impact the reported results (Park et al., 2024). To ensure fairness, we use a widely adopted test set focused on frontal-view CXRs, aligned with previous studies such as MAIRA-1 (Hyland et al., 2024) and LLaVA-Rad (Chaves et al., 2024) 444The test set consists of 2,461 frontal-view image-report samples..
Additionally, the latest concurrent work M4CXR (Park et al., 2024) employs multi-turn chain-of-thought (Wei et al., 2023) prompting to generate reports, which differs from our task setup. Furthermore, we do not compare with the recent work MAIRA-2, as it is a MLLM specifically designed for grounded radiology report generation task, which incorporates lateral views and prior study reports for each subject within the input prompt. Bannur et al. (2024) emphasises a positive transfer between this task and general report generation, which is also beyond the scope of our study 555For comparisons with the latest concurrent and non-LLM-based models can be found in Appx. D..
Taking these considerations into account, we compared our model with state-of-the-art models, including LLaVA-Med (Li et al., 2023a), CheXagent (Chen et al., 2024), GPT-4V (OpenAI et al., 2024), Med-PaLM (Tu et al., 2023), LLaVA-Rad and MAIRA-1. The results are shown in Table 2. Since many of these models are not publicly available, we present their evaluation results as reported in the original sources.
| Lexical: | |||||||
| ROUGE-L | 27.6 | 21.5 | 13.2 | 27.5 | 30.6 | 28.9 | 36.7 (19.9%) |
| BLEU-1 | 35.4 | 16.9 | 16.4 | 32.3 | 38.1 | 39.2 | 51.3 (30.9%) |
| BLEU-4 | 14.9 | 4.7 | 17.8 | 11.5 | 15.4 | 14.2 | 24.5 (37.6%) |
| METEOR | 35.3 | – | – | – | – | 33.3 | 48.9 (38.5%) |
| Clinical: | |||||||
| RadGraph-F1 | 19.1 | – | – | 26.7 | – | 24.3 | 32.9 (23.2%) |
| RG | 23.8 | 20.5 | 13.2 | – | 29.4 | 29.6 | 37.6 (27.0%) |
| RadCliQ() | 3.3 | – | – | – | – | 3.1 | 2.7 (12.9%) |
| CheXbert vector | 36.9 | – | – | – | – | 44.0 | 46.9 (6.59%) |
| CheXpert-F1: | |||||||
| Micro-F1-14 | 42.7 | 39.3 | 35.5 | 53.6 | 57.3 | 55.7 | 55.9 (-2.4%) |
| Macro-F1-14 | 26.9 | 24.7 | 20.4 | 39.8 | 39.5 | 38.6 | 40.4 (1.5%) |
| Micro-F1-5 | 43.9 | 41.2 | 25.8 | 57.9 | 57.4 | 56.0 | 60.1 (3.8%) |
| Macro-F1-5 | 36.3 | 34.5 | 19.6 | 51.6 | 47.7 | 47.7 | 53.8 (4.3%) |
From Table 2 we can see that, for all traditional lexical metrics, Libra substantially outperforms the baseline models, achieving scores of 36.7 for ROUGE-L, 51.3 and 24.5 for BLEU-1 and BLEU-4 respectively, and 48.9 for METEOR. In clinical metrics, Libra also showed outstanding performance, substantially improving in RadGraph-F1 and its variant RG, with scores of 32.9 and 37.6, respectively. Notably, Libra achieved the highest score (i.e. 2.7) on the radiologist-aligned RadCliQ metric. For CheXbert-derived metrics, it also secured the top position with a vector embedding matrix score of 46.9. In the CheXpert 5-class subset, Libra excelled with Micro-F1 and Macro-F1 scores of 60.1 and 53.8, respectively. In the CheXpert 14-class subset, Libra achieved the highest Macro-F1 score of 40.4, while its Micro-F1 score of 55.9 was only slightly below the leading model.
In summary, Libra demonstrated outstanding performance across both sets of metrics, with only minor gaps in one specific measure. The remarkable capabilities of Libra can be attributed to its innovative Temporal Alignment Connector which effectively captures temporal changes in the dataset, enabling the generation of highly accurate and clinically relevant reports.
4 Ablation Study
To thoroughly assess the impact of various optimisation components, we conducted a series of ablation experiments, evaluating different modules and dataset expansions within Libra. All ablation models were tested on the MIMIC-CXR official test split for the Findings section generation task, using identical hyperparameter settings during training and inference.
Q1: Does incorporating temporal information bring positive effects to Libra in RRG tasks?
| Metric | Libra | |
| w/ true | w/ dummy(%) | |
| Lexical: | ||
| ROUGE-L | 36.66 | 36.17 (-1.34%) |
| BLEU-1 | 51.25 | 51.20 (-0.10%) |
| BLEU-4 | 24.54 | 24.33 (-0.86%) |
| METEOR | 48.90 | 48.69 (-0.43%) |
| BERTScore | 62.50 | 61.94 (-0.90%) |
| F1 | 35.34 | 32.72 (-7.41%) |
| Clinical: | ||
| RadGraph-F1 | 32.87 | 32.42 (-1.37%) |
| RG | 37.57 | 36.92 (-1.73%) |
| RadCliQ() | 2.72 | 2.76 (-1.47%) |
| CheXbert vector | 46.85 | 46.31 (-1.15%) |
| CheXpert-F1: | ||
| Micro-F1-14 | 55.87 | 55.25 (-1.11%) |
| Macro-F1-14 | 40.38 | 40.15 (-0.57%) |
| Micro-F1-5 | 60.07 | 58.93 (-1.90%) |
| Macro-F1-5 | 53.75 | 52.61 (-2.12%) |
Temporal information is embedded in paired images and referenced in the corresponding radiology reports, reflecting changes over time through references to previous symptoms and descriptions of their progression. As indicated in Table 1, 86% of the test data includes true prior images. To assess whether Libra can effectively perceive and utilise temporal information during inference, we conducted separate evaluations using either true prior images or only dummy prior images. When true prior images were unavailable, the model treated the current image as a dummy prior image. As shown in Table 3, Libra demonstrated substantial enhancements across all metrics when true prior images were used as references. Specifically, the improvement in clinical scores exceeds that of lexical scores, indicating that temporal information is crucial for generating high-quality medical reports, beyond merely improving linguistic fluency. Notably, the F1 score shows the most significant impact, with a difference of 7.41%. This highlights Libra’s capability to effectively leverage temporal changes provided by true prior images, thereby enhancing the quality of the generated Findings section and overall performance in the RRG task.
Q2: Does the Temporal Alignment Connector effectively improve model performance?
| Metric | Libra-1 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
| Lexical: | |||||
| Clinical: | |||||
| CheXpert-F1: | |||||
To evaluate the impact of TAC on Libra’s performance in the RRG task, we initialised a model with the RAD-DINO (Pérez-García et al., 2024) image encoder, TAC, and Meditron (Chen et al., 2023b) as the LLM. A baseline experiment (Libra-1) was conducted by fine-tuning only the TAC for the Findings generation task. As shown in Table 4, we performed ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC. Removing TFM restricted the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023) but with a four-layer MLP to align the image representation with the LLM’s hidden dimensions. Without LFE, the model follows the LLaVA setup, utilising the penultimate layer of the image encoder. The results indicate that removing any TAC submodule leads to a decline in all metrics compared to Libra-1. Removing TFM caused a notable drop in the F1 score (2%), emphasising its role in capturing temporal information. Removing LFE especially decreases RadGraph-related scores, highlighting its importance in extracting detailed image features. The removal of PIPB impacted clinical scores more than lexical scores. Finally, removing the entire TAC results in substantial declines across all metrics, underscoring its critical role in integrating image details and temporal information666For more ablation studies, including the impact of radiology-specific pre-trained models, see Appx. C..
Q3: Are additional Impression and VQA datasets needed during the feature alignment?
| Metric | Stage: 1 | Stage: 2 | ||
| Libra | Libra-f | Libra | Libra-f | |
| Lexical: | ||||
| ROUGE-L | 27.27 | 27.56△ | 36.66 | 35.31▼ |
| BLEU-1 | 41.24 | 34.84▼ | 51.25 | 49.92▼ |
| BLEU-4 | 13.59 | 11.51▼ | 24.54 | 23.05▼ |
| METEOR | 39.44 | 35.50▼ | 48.90 | 47.99▼ |
| BERTScore | 56.00 | 55.87▼ | 62.50 | 61.28▼ |
| F1 | 24.80 | 26.63△ | 35.34 | 33.52▼ |
| Clinical: | ||||
| RadGraph-F1 | 20.45 | 22.52△ | 32.87 | 30.77▼ |
| RG | 25.19 | 27.32△ | 37.27 | 35.44▼ |
| RadCliQ () | 3.31 | 3.10△ | 2.72 | 2.83▼ |
| CheXbert vector | 35.33 | 42.02△ | 46.85 | 45.32▼ |
| CheXpert-F1: | ||||
| Micro-F1-14 | 43.63 | 52.48△ | 55.87 | 54.11▼ |
| Macro-F1-14 | 25.68 | 36.87△ | 40.38 | 37.16▼ |
| Micro-F1-5 | 49.75 | 56.63△ | 60.07 | 58.76▼ |
| Macro-F1-5 | 40.40 | 49.33△ | 53.75 | 51.99▼ |
To evaluate the impact of dataset expansion during the first stage on Libra’s performance, we trained a model (Libra-f) using only the Findings data in the first stage, while Libra incorporated additional Impression and VQA data for alignment, in Table 5. After the first stage, Libra showed superior performance in lexical metrics compared to Libra-f, but a slight decline in clinical metrics. This can be attributed to the inclusion of VQA tasks, which encouraged the model to focus on more fine-grained and grounded information. While VQA emphasises the detailed description of the single symptom, the Findings section generation task requires a comprehensive overview of the CXR, encompassing multiple normal and abnormal findings. This shift in focus also impacted the F1 score, as each finding typically involves its own temporal changes. Additionally, the reduction in the number of generated findings entities led to a decrease in the number of identified temporal entities. In the second stage, fine-tuned on the Findings dataset, Libra achieved the best results across all metrics, indicating that incorporating additional datasets in the first stage enhances Libra’s understanding and reasoning of CXRs.
5 Performance Analysis
We used CXRs from the official test split to generate Findings sections and perform a qualitative analysis of Libra. Figure 2 illustrates a case where no true prior image was available for reference. For example, while the ground truth report only mentioned the presence of “sternal wires,” Libra not only identified their presence but also provided details on the specific type. This demonstrates Libra can deliver more precise clinical information by not only recognizing conditions but also distinguishing the specific types of surgical interventions.
Original
report
The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Sternal wires are intact.
Libra
The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Median sternotomy wires are intact.
For the samples with true prior images, Libra accurately describes the presence and changes in radiological conditions. As shown in the Figure 3, the current image revealed additional findings of pleural effusion and pneumonia compared to the prior image. The model effectively detected these progressive changes, provided detailed descriptions, and suggested potential causes for further investigation. Notably, it explicitly mentioned the comparative information relative to the prior image, offering a clear understanding of the temporal changes.
Current image
Prior image
Original report (current image)
In comparison
with the study of , there is again evidence of a superior mediastinal mass consistent with the known thyroid enlargement that is better seen on the recent CT scan. Continued bibasilar opacification most likely represents atelectasis, though in the appropriate clinical setting, superimposed pneumonia would have to be considered. The cardiac silhouette is within normal limits, though there may be some mild elevation of pulmonary venous pressure. Small pleural effusions are probably present bilaterally.
Libra
As compared to
radiograph, cardiomediastinal contours are stable. New patchy and linear opacities have developed at the lung bases, right greater than left, and may reflect patchy atelectasis, aspiration, or developing infectious pneumonia. Small pleural effusions are present bilaterally.
w/ dummy prior image
Cardiac silhouette is upper limits of normal in size and accompanied by pulmonary vascular congestion and minimal interstitial edema. Patchy bibasilar opacities are present, and could reflect patchy atelectasis, aspiration, or developing infectious pneumonia. Followup radiographs may be helpful in this regard. Small bilateral pleural effusions are also demonstrated.
Swap order
Cardiomediastinal contours are normal. Previously present bibasilar opacities have nearly completely resolved with minimal residual linear opacification remaining at the right lung base. No new areas of consolidation are identified, and there are no pleural effusions.
Original report (prior image)
The heart is normal in size. The mediastinal and hilar contours appear stable. The aortic arch is calcified. There is no definite pleural effusion or pneumothorax. Streaky opacity in the right costophrenic angle suggests minor atelectasis, which has decreased. There is similar mild spinal curvature.
To further assess the model’s robustness, we conducted a more detailed analysis of this case. First, we generated a report using only the dummy prior image as a reference. While Libra successfully detected abnormal findings, it failed to specify changes relative to the previous image, describing only the current state without indicating disease progression. Moreover, it provided less clarity in distinguishing the distribution and severity of abnormalities. This suggests that, without a true prior image for comparison, the model struggles to accurately describe the progression of findings.
Evaluating Temporal Consistency in Libra’s Report Generation To investigate the model’s ability to capture temporal changes, we swapped the order of the images, using the true prior image as the current image and vice versa. Interestingly, the generated report indicated an improvement in the patient’s condition, with most abnormalities resolved and no new findings detected, which is contrary to the ground truth of the original current image report. Moreover, the report generated from the swapped images closely resembled the original report of the true prior image, as shown at the bottom of Figure 3, which described a stable or improving condition for the patient. This striking observation demonstrates that Libra can effectively utilise both current and prior studies to generate accurate reports for assessing changes, simulating the conditions of standard clinical practice.
6 Conclusion
In this study, we presented Libra, a temporal-aware multimodal large language model specifically designed for chest X-ray report generation tasks. Trained exclusively on the open-access MIMIC-CXR dataset (Johnson et al., 2019b), Libra utilises a two-stage training framework that leverages a radiology-specific pre-trained image encoder and language model, connected through a specially designed Temporal Alignment Connector to bridge visual and textual modalities. By identifying observations from current CXR and referencing prior scan, Libra generates radiology reports with substantial improvements across all metrics compared to existing models of the same parameter scale. Qualitative analysis in the case study highlights Libra’s ability to effectively utilise the temporal relationships between current and prior images, addressing the challenge of hallucinations related to prior study references in RRG tasks. Libra sets a new paradigm for MLLMs in multimodal medical AI research. In the future, we aim to further enhance Libra’s clinical applicability and accuracy, paving the way for the development of general-purpose AI models in the medical domain. For a detailed discussion of Libra’s limitations and future work, please refer to Appx. A.3.
Appendix A Appendix A
A.1 Related Work
Radiology Report Generation Several efforts have been made to develop models tailored for radiology report generation (RRG), driven by the need to address the long-tail distribution of observations in chest X-rays and provide fine-grained descriptions of findings. This task has gained prominence as an important objective for the designed system (Wang et al., 2018b).
Such systems inherently require a language generation component. Initially, recurrent neural networks were used (Liu et al., 2019), but these have recently been replaced by Transformer architectures (Miura et al., 2021b; Chen et al., 2022), including large language models (LLMs) such as PaLM (Chowdhery et al., 2022) and Vicuna-7B (Chiang et al., 2023).
To ensure clinical accuracy in generated text, some studies have moved beyond standard language modelling losses, instead employing reinforcement learning (RL) to optimise for rewards that prioritise “clinically relevant” features, such as the inclusion of specific findings (Liu et al., 2019; Irvin et al., 2019b) or maintaining logical consistency (Miura et al., 2021a; Delbrouck et al., 2022a). However, these methods often depend on external models like CheXbert (Smit et al., 2020) or RadGraph (Jain et al., 2021) to extract clinical entities, adding complexity to the optimisation process.
With the advancement of LLMs, an increasing number of studies have demonstrated that plain auto-regressive language modelling can achieve substantial results. However, we acknowledge that the benefits of more complex training objectives or RL-based approaches may be complementary. Consequently, research on leveraging temporal information in RRG tasks can be broadly categorised into two types: LLM-based methods and non-LLM-based methods.
non-LLM-based Model These architectures are typically transformer encoder-decoder models and their variants. They are often composed of multiple modules that require separate training. Meanwhile, they handle single and double image inputs by symbolically differentiating tasks and employing two distinct neural network architectures: one designed for single-image inputs and another for double-image inputs.
Serra et al. (2023) utilises symbolic alignment in its Longitudinal Projection Module and employs a separately trained BERT-based (Devlin et al., 2019) text generator. RECAP (Hou et al., 2023a) implements a transformer encoder-decoder with symbolic task differentiation and two-stage training, beginning with classification tasks followed by report generation. TiBiX (Sanjeev et al., 2024) adopts a transformer model with causal attention layers and incorporates learnable padding tokens to handle cases without prior images. BioViL-T (Bannur et al., 2023) is a self-supervised vision-language training framework that features a CNN–Transformer hybrid multi-image encoder trained jointly with a BERT-based text model.
On one hand, the difference in model parameter sizes, and on the other, as LLM-based models generally outperform other types of models in the RRG task, papers on non-LLM-based models or those using small language models typically do not compare their methods with LLM-based approaches. At the same time, we conducted comparisons and discussions to reaffirm this point, as detailed in Appx. D.2.
LLM-based Model Current LLM-based models have achieved notable success in the RRG task, largely benefiting from visual instruction tuning (Liu et al., 2023). Structurally, these models (Li et al., 2023a; Chaves et al., 2024; Hyland et al., 2024; Zhou et al., 2024; Park et al., 2024) typically consist of an image encoder and an adapter that bridges the connection to the LLM. These adapters often rely on single-layer hidden representations (e.g., the last or penultimate layer) from pre-trained image encoders, limiting their ability to simultaneously integrate information from multiple images.
In end-to-end training, LLM-based models handle multiple image inputs by concatenating them with textual prompts to form a composite input for the LLM. For example, the input format is “<Current Image Placeholder> + <Prior Imgae Placeholder> + <Prompt>”.
Although our model adopts an end-to-end training paradigm, it does not rely on the vanilla approach to handle multiple images. Instead, we designed a novel adapter called the Temporal Alignment Connector (TAC). On one hand, TAC leverages all hidden layer features from the image encoder, setting it apart from the default configurations used in previous works. On the other hand, TAC models temporal information directly rather than simply concatenating two images in the latent space of the LLM.
Radiological Image Representation As radiology lies outside the distribution of most general-domain image models, a radiology-specific pre-trained model is crucial for the RRG task (Pérez-García et al., 2024).
Zhou et al. (2023) introduces a unified framework, Masked Record Modeling (MRM), which combines self-supervision and radiology report supervision to enhance radiograph representation learning. BioViL-T (Bannur et al., 2023) integrates a CNN-Transformer hybrid architecture for multimodal modelling, leveraging temporal structures in medical imaging and reports to perform tasks such as disease progression classification and report generation. BiomedCLIP (Zhang et al., 2024b) is a multimodal biomedical foundational model pre-trained on various biomedical tasks. RAD-DINO (Pérez-García et al., 2024) is a medical image encoder that employs a pure image-based self-supervised learning approach from DINOv2 (Oquab et al., 2024) for continuous pretraining, focusing exclusively on image data to avoid the limitations of text supervision.
In recent parallel works, RAD-DINO has been widely applied to RRG tasks, including MAIRA-2 (Bannur et al., 2024), M4CXR (Park et al., 2024). Furthermore, Pérez-García et al. (2024) highlighted that RAD-DINO outperforms other types of image encoders in RRG tasks. Therefore, our model also adopts RAD-DINO as the image encoder.
A.2 Research Objectives
Temporal Information Temporal change is a crucial factor in understanding disease progression, especially in radiology, where paired images and corresponding radiology reports capture changes over time. This temporal context is often reflected in comparisons between current and prior scans, documenting symptom evolution or newly identified findings.
The relativity of position within the timeline is a key determinant of temporal information. This relative positioning does not alter the established description of well-established facts about symptoms already present in the current image. For instance, if the prior image is recent, descriptions of changes will be minimal, whereas an older prior image may highlight more pronounced changes.
Research Object Our study focuses on the chest X-ray (frontal view), rather than the individual patient. Specifically, we aim to generate a report for the current image while referencing the temporal information provided by a single prior image as auxiliary information.
Our model is primarily designed to handle scenarios with limited temporal information in the RRG task. For instance, in a case where a patient has two scans taken just milliseconds apart, the current and prior images would be nearly identical, as no pathological changes would manifest within such a short interval. This extreme scenario demonstrates how the model handles clinical practice under limited temporal information. In such cases, the correct diagnosis for this minimal interval would be that the patient’s condition is “stable.” In our design, the dummy prior image addresses scenarios where a true prior image is absent.
However, patients often undergo multiple prior scans, which may include temporal information from different orientations. This more complex temporal information lies beyond the scope of our current study. We will provide a detailed discussion of this scenario in Appx. A.3.
A.3 Limitations and Future Work
Handling Multiple Prior Scans Our current model is designed to process a single prior scan alongside the current scan. While this approach aligns with standard clinical workflows, which typically prioritise the most recent prior study for comparisons, it overlooks scenarios where multiple prior scans could offer a richer temporal perspective. For instance, analysing a sequence of images spanning an extended period could provide deeper insights into gradual disease progression. Future efforts should focus on extending our framework to incorporate multiple prior scans efficiently, enabling a more nuanced understanding of temporal patterns in clinical data.
Temporal Information Beyond Image Comparisons At present, our model captures temporal information through paired image comparisons and their corresponding textual reports. However, clinical assessments often draw upon a broader context, including historical notes, laboratory results, and other longitudinal patient data. Expanding our approach to integrate these diverse temporal data sources could facilitate a more holistic understanding of disease trajectories and patient history, significantly enhancing clinical applicability.
Challenges with Sparse Temporal Data In cases where prior scans are unavailable or only minimally informative (e.g., scans taken within a very short time interval), our use of a “dummy prior image” addresses the lack of temporal data. However, the model’s ability to interpret and generate meaningful outputs under these constraints may still be limited. Future work could explore methods for imputing or synthesising temporal context to address such cases more effectively.
Computational Complexity Our use of temporal alignment mechanisms and multi-layer feature integration introduces increased computational demands. This complexity poses challenges for real-time deployment, particularly in resource-constrained clinical environments.
Generalisability Across Modalities and Datasets This study focuses exclusively on frontal-view chest X-rays and utilises the MIMIC-CXR dataset (Johnson et al., 2019b). However, our approach has not been evaluated on other imaging modalities (e.g., CT or MRI) or datasets such as CheXpert (Irvin et al., 2019b) or PadChest (Bustos et al., 2020).
Based on the identified limitations, we propose several directions for future research and development:
Develop frameworks for integrating multiple prior scans with dynamic temporal reasoning.
Extend the model to handle multi-modal imaging and textual data for comprehensive diagnostics.
Investigating the integration of temporal data from diverse clinical sources, including electronic health records (EHRs)
These enhancements will not only address the current limitations but also extend the potential of temporal-aware multimodal models in radiology and beyond.
Appendix B Appendix B
B.1 Training Configuration
We train the model using a standard auto-regressive language modelling loss (cross-entropy). For this study, we utilise Meditron-7b (Chen et al., 2023b) as the LLM, with a total batch size of 16 throughout the training process. The training is conducted on a computational infrastructure equipped with A6000 GPU (48GB of memory) and configured with DeepSpeed ZeRO-3 optimization (Rajbhandari et al., 2020) and BF16 precision enabled. A cosine learning rate scheduler is employed, starting with a warm-up phase of . In the first stage of training, we run for 1 epoch with a learning rate of . In the second stage, the model is trained for 3 epochs at the same learning rate. The LoRA (Hu et al., 2021) parameters are set to and . The final checkpoint for all runs is selected based on the observation of the minimum loss on the evaluation dataset throughout the training process.
B.2 Detailed Description of Datasets
MIMIC-CXR. (Johnson et al., 2019b) This is a large, publicly accessible dataset comprising 377,110 DICOM images across 227,835 studies, each accompanied by a radiology report (Johnson et al., 2019b). For images, we use the commonly available JPEG files from MIMIC-CXR-JPG (Johnson et al., 2019a), rather than the original DICOM files, and we preprocess the dataset to exclude non-AP/PA scans. For each report, we extract the Findings, Impression, Indication, History, Comparison, and Technique sections using rule-based heuristics supported by the official MIMIC code repository (Johnson et al., 2018).
For the Findings section generation task, we discard studies where the Findings could not be extracted, while allowing for missing content in other clinical sections. The same approach is applied to the Impression section generation task. In all our experiments, we adhere to the official MIMIC-CXR dataset split. Additionally, we retrieve prior images by following the chronological order of studies as indicated by the official labels, selecting the closest prior study as the reference image. It is important to note that, to prevent data leakage between the train, validation, and test sets, prior images are retrieved only from within the same split.
Medical-Diff-VQA. (Hu et al., 2023) This dataset is a derivative of the MIMIC-CXR dataset, focused on identifying differences between pairs of main and reference images. The data split adheres to the original labelling, ensuring no data leakage occurs. In total, this dataset comprises 700,703 question-answer pairs derived from 164,324 main-reference image pairs. The questions are divided into seven categories: abnormality, location, type, view, presence, and difference. Each pair consists of a main (current) image and a reference (prior) image, both taken from different studies of the same patient. The reference image is always selected from an earlier visit, with the main image representing the later visit. Of the seven question types, the first six types focus on the main image, while the “difference” questions involve both images.
MIMIC-Ext-MIMIC-CXR-VQA. (Bae et al., 2023) This dataset is an extension of the MIMIC-CXR dataset, specifically designed for VQA tasks within the medical domain, specifically focused on CXRs. It includes questions generated from 48 unique templates covering seven content types: presence, anatomy, attribute, abnormality, size, plane, and gender. Each template was developed with the guidance of board-certified medical experts to ensure clinical relevance, addressing both standard medical VQA content and more complex logical scenarios. In total, the dataset consists of 377,391 unique entries. Since annotations are based on single images, the current image serves as a dummy prior image for all entries in our experiment.
For this study, we carefully selected datasets that provide complete reports and temporal information (i.e., prior images) to ensure alignment with our research objectives for the RRG task. After thoroughly evaluating other datasets, we found them unsuitable for the following reasons:
CheXpert. (Irvin et al., 2019b) This dataset includes annotated scans with label-specific annotations rather than full medical reports. While useful for training image encoders or annotation models, it is not appropriate for the RRG task, which requires complete diagnostic reports.
PadChest. (Bustos et al., 2020) Although it includes reports and corresponding prior images, its reports are in Spanish, placing cross-language training beyond the scope of our model.
IU-Xray. (Demner-Fushman et al., 2016) This dataset lacks patient-level metadata and prior study information, which is critical for our focus on temporal information in chest X-rays.
Chest ImaGenome Dataset. (Wu et al., 2021a)(a derivative of MIMIC-CXR) This dataset does not follow the official MIMIC-CXR split, risking data leakage across train, validation, and test sets.
B.3 Evaluation Metrics Details
[…]

### v2

@@ Appendix E Datasets Description
MIMIC-CXR
(Johnson et al., 2019b) This is a large, publicly accessible dataset comprising 377,110 DICOM images across 227,835 studies, each accompanied by a radiology report (Johnson et al., 2019b). For images, we use the commonly available JPEG files from MIMIC-CXR-JPG (Johnson et al., 2019a), rather than the original DICOM files, and we preprocess the dataset to exclude non-AP/PA scans. For each report, we extract the Findings, Impression, Indication, History, Comparison, and Technique sections using rule-based heuristics supported by the official MIMIC code repository (Johnson et al., 2018).
For the Findings section generation task, studies without extractable Findings are discarded, while other missing sections are permitted. The same approach is applied to the Impression section generation task. In all our experiments, we adhere to the official MIMIC-CXR dataset split.
Meanwhile, we retrieve prior images by following the chronological order of studies as indicated by the official labels, selecting the closest prior study as the reference image. It is important to note that, to prevent data leakage between the train, validation, and test sets, prior images are retrieved only from within the same split.
Medical-Diff-VQA
(Hu et al., 2023) This dataset is a derivative of the MIMIC-CXR dataset, focused on identifying differences between pairs of main and reference images. The data split adheres to the original labelling, ensuring no data leakage occurs. In total, this dataset comprises 700,703 question-answer pairs derived from 164,324 main-reference image pairs. As shown in Table 5, the questions are divided into seven categories: abnormality, location, type, view, presence, and difference.
Each pair consists of a main (current) image and a reference (prior) image, both taken from different studies of the same patient. The reference image is always selected from an earlier visit, with the main image representing the later visit. Of the seven question types, the first six types focus on the main image, while the “difference” questions involve both images.
MIMIC-Ext-MIMIC-CXR-VQA
(Bae et al., 2023) This dataset extends MIMIC-CXR for VQA tasks tailored to CXRs. It includes questions generated from 48 unique templates covering seven content types: presence, anatomy, attribute, abnormality, size, plane, and gender, as shown in Table 5. Each template was developed with the guidance of board-certified medical experts to ensure clinical relevance, addressing both standard medical VQA content and more complex logical scenarios. In total, the dataset consists of 377,391 unique entries. Since annotations are based on single images, the current image serves as a dummy prior image for all entries in our experiment.
For this study, we carefully selected datasets that provide complete reports and temporal information (i.e., prior images) to ensure alignment with our research objectives (see Appx. B) for the RRG task. After thoroughly evaluating other datasets, we found them unsuitable for the following reasons:
CheXpert
(Irvin et al., 2019) This dataset includes annotated scans with label-specific annotations rather than full medical reports. While useful for training image encoders or annotation models, it is not appropriate for the RRG task, which requires complete diagnostic reports.
PadChest
(Bustos et al., 2020) Although it includes reports and corresponding prior images, its reports are in Spanish, placing cross-language training beyond the scope of our model.
IU-Xray
(Demner-Fushman et al., 2016) This dataset lacks patient-level metadata and prior study information, which is critical for our focus on temporal information in chest X-rays.
Chest ImaGenome Dataset
(Wu et al., 2021) Although derived from MIMIC-CXR (Johnson et al., 2019b), it does not follow the official split, raising concerns about potential data leakage between training, validation, and test sets.
Meanwhile, the following two datasets were processed using GPT-4 (OpenAI et al., 2024) to eliminate hallucinated references to prior exams. While this prevents erroneous comparisons, it also removes essential temporal information originally present in the reports, potentially affecting tasks that rely on temporal reasoning.
LLaVA-Rad MIMIC-CXR Dataset
(Chaves et al., 2024) This dataset was refined using GPT-4 (OpenAI et al., 2024) through a structured text-cleaning pipeline. The process involved: (1) correcting typographical errors and split words, (2) removing redundant or repeated phrases to improve clarity, (3) eliminating explicit temporal references (e.g., “Compared to the prior study, no significant interval change was noted”) to ensure the report focuses exclusively on the current image, and (4) restructuring content into standardised sections, including Indication, Findings, and Impression.
ReXPref-Prior Dataset
(Banerjee et al., 2024) A modified version of MIMIC-CXR (Johnson et al., 2019b) in which GPT-4 (OpenAI et al., 2024) systematically removes all references to prior exams from both the Findings and Impression sections. While this adjustment prevents spurious prior-study references, it also eliminates crucial temporal context, limiting its suitability for applications requiring longitudinal assessment of disease progression.
Appendix F Evaluation Metrics
F.1 Lexical Metrics
We employed standard natural language generation metrics to quantify the overlap between generated and reference reports. Specifically, ROUGE-L (Lin, 2004) measures the length of the longest common subsequence between the generated and reference reports. BLEU- (Papineni et al., 2002) calculates n-gram precision and applies a brevity penalty to discourage overly short predictions. METEOR (Banerjee and Lavie, 2005), computes the weighted harmonic mean of unigram precision and recall, with an additional penalty for fragmenting consecutive word sequences. Finally, we report BERTScore (Zhang et al., 2020a), which leverages pre-trained contextual embeddings from BERT (Devlin et al., 2019) to match words in candidate and reference sentences based on cosine similarity. We used default parameters for all of these evaluation metrics.
[…]